FOURIER TRANSFORM INFRARED SPECTROSCOPY OF CH_{3}OO RADICAL IN MID-INFRARED RANGE

KUO-HSIANG HSU and YUAN-PERN LEE, Department of Applied Chemistry and Institute of Molecular Science, National Chiao Tung University, Hsinchu 30010, Taiwan and Institute of Atomic and Molecular Sciences, Academia Sinica, Taipei 10617; \underline{MENG HUANG} and TERRY A. MILLER, Laser Spectroscopy Facility, Department of Chemistry, The Ohio State University, 120 W. 18th Avenue, Columbus, Ohio 43210.

A mid-infrared spectrum of the CH_{3}OO radical at room temperature has been measured by Fourier-transform infrared spectroscopy. The CH_{3}OO radicals were produced by photolysis of a CH_{3}I/O_{2} mixture at 248 nm or a CH3COCH3/O2 mixture at 193 nm; the total pressure is 100 Torr and the precursor is about 1.6-2.0%. The ν_{2}, ν_{5}, ν_{6}, ν_{7}, and ν_{9} fundamental bands with origins at 2954.0, 1182.6, 1118.0, 910.8, and 3021.4 cm^{−1} have been observed, which are in good agreement with previous low-resolution work.^{a} Particular attention has been given to simulate the rotational structure of the ν_{2} band. Sequence band structure from the methyl torsion mode ν_{12} was included in the simulation of this band as well as some transitions from the precursor. The simulation shows generally good consistency with the experimental spectrum and allows the determination of the molecule’s rotational constants.

aD.-R. Huang, L.-K. Chu, and Y.-P. Lee J. Chem. Phys. 127, 7 (2007)